OpenStudy (anonymous):

How to use leibniz notation to solve y=sqr 2u , u=6-x x=-3

OpenStudy (anonymous):

Wait is the sqr representing a square-root symbol?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

i did the question. I got (u^-1/2)(2)(-1)

OpenStudy (anonymous):

So it's: \[y=\sqrt{2u}, u=6-x,x=-3\]Right? @0213

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The result would be a constant then, so the answer would be 0, correct?

OpenStudy (anonymous):

the answer at the back of my book is 1/2sqr3

OpenStudy (anonymous):

sorry its -1/3sqr2

OpenStudy (anonymous):

So first, replace u with 6-x in the square root:\[y=\sqrt{2(6-x)} \rightarrow y=\sqrt{12-2x}\]Now just differentiate with power rule and chain rule.\[y'=\frac{ 1 }{ 2 }(12-2x)^{-\frac{ 1 }{ 2 }}*-2 \rightarrow y'=-\frac{ 2 }{ 2\sqrt{12-2x} } \rightarrow y'=-\frac{ 1 }{ \sqrt{12-2x} }\]

OpenStudy (anonymous):

And the derivative here at x = -3 is then:\[y'(-3)=-\frac{ 1 }{ \sqrt{12-2(-3)} }\rightarrow y'(-3)=-\frac{ 1 }{ \sqrt{12+6} } \rightarrow y'(3)=-\frac{ 1 }{ \sqrt{18} }\]

OpenStudy (anonymous):

@0213

OpenStudy (anonymous):

but i have to use leibniz notation. I have to show the derivative of y multplied by the derivative of u . Thats the way I'm doing it .

OpenStudy (anonymous):

Oh you want to show it in terms of du/dx?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Should've said that before -.- \[y'=\frac{ 1 }{ 2 }(2u)^{-\frac{ 1 }{ 2 }}*2*\frac{ du }{ dx } \rightarrow y'=\frac{ 1 }{ \sqrt{2u} }*\frac{ du }{ dx } \rightarrow y'=\frac{ 1 }{ \sqrt{12-2x} }*\frac{ du }{ dx }\]@0213

OpenStudy (anonymous):

so I won't multiply the 2 and the 1/2 when I bring it down in the first part of the equation

OpenStudy (anonymous):

First part is power rule and chain rule. You take the power of 1/2 to the front of 2u and the power of 2u becomes -1/2, that's the power rule. Now you have to differentiate the stuff under the square root as well and multiply it by that. This is the chain rule. So when you differentiate 2u with constant multiple rule, you get 2*du/dx. So now you multiply the whole thing by 2*du/dx but the 2 in the denominator of 1/2 and the 2 being multiplied by du/dx cancel out. From there you just simplify and in the last step, I just replaced the u in the square root with 6 - x to make it the root of 12 - 2x. @0213

OpenStudy (anonymous):

so when I have a square root with a value like 3x or 5x-1 , I have to look at the square root alone and take the derivative of it and then look at the value inside and take the derivative of it